# Steady-states and response functions of the periodically driven O($N$) scalar field theory

### v1

@@ III.1.4 Effect of finite temperature: Fluctuations shift the critical point
In the large frequency limit, $\Omega\gg\sqrt{|r_{1}|}$, $\gamma$, we can also approximate $C_{n}(q)\approx\frac{T}{q^{2}+\lambda_{0}}\delta_{n,0}$ and then calculate the critical value of $r_{0}$ for which the superconducting phase emerges. To this end, we use the definiton of $\lambda_{0}$ as in Eq. (9) at $n=0$ and setting $b_{n}=0$. Then, we make use of the stability condition Eq. (11) (and replace $r_{0}$ with $\lambda_{0}$). The new critical value for $r_{0,c}$ is then given by:
@@ IV Full numerical solution
Finite-momentum instabilities— We also performed a stability analysis of each phase obtained by numerically solving Eqs. (4). We found that especially for larger values of $r_{1}/\Omega^{2}$ the normal phase is unstable to a finite-momentum symmetry broken state, where the order parameter displays oscillations as a function of time with a characteristic frequency of $\Omega$ (period-preserving state) of $\Omega/2$ (period-doubling state). For details of this analysis see App. D. As shown in App. E, such phases also exhibit a Meissner effect, which is why we dubbed them as pair density waves (PDW) in Fig. 1 (marked by red or blue hatches). For a finite-momentum instability, the order parameter can take two different forms:
@@
with $f_{\nu}(t)$ a time-periodic function with period $(2\pi)/\nu$ and $\mathbf{Q}$ a particular wave vector where the instability occurs. In the first case the O($N$) symmetry is broken down to O(N1N-1) and both the translation and the spatial rotation symmetries are broken. In the second case, a subgroup of the combination of translations and O($N$) rotations remains preserved, while spatial rotations are broken and the global O($N$) symmetry is broken down to O(N2N-2). More complex patterns are in principle possible as one is allowed to construct any linear combination of wave vectors $\mathbf{q}$ satisfying $|\mathbf{q}|=|\mathbf{Q}|$. The determination of which finite-$q$ configuration is realized is beyond our scope and it would require solving a generalization of Eqs. (4) where $\bar{b}$ becomes spatially nonuniform. We also remark that our analysis (detailed in App. D) relies on the assumption that the finite-$q$ order parameter develops continuously upon crossing the transition line. If the transition happens to be first-order, the boundaries of the finite-momentum phases in Fig. 1 might change. We also found a regime (yellow region in Fig. 1), where $\lambda(t)$ is not a periodic function. Here, the correlation function $C_{\mathbf{q}}(t)$ has a large peak at finite $\mathbf{q}$, which is why we dubbed this phase a quasi-periodic PDW. Finite momentum instabilities in parametrically driven bosonic systems were discussed also in Refs. [Dai2021; Chandra25a; Chandra25b; Kaplan2025b].
@@ V.1 Meissner effect
Figure 5: Behavior of the magnetic field inside the material for a superconducting state with an order parameter that is either uniform or carries a finite momentum in the form of Eq. (17b). The $x$-axis coordinate indicates the distance from an interface to vacuum. In panels (a) and (b) we show a state with period doubling, whereas in panels (c) and (d) we have a superconducting state without period doubling, with an order parameter that oscillates around a finite offset. Panels (a) and (c) display the magnetic field profiles near the vacuum interface, in contrast to panels (b) and (d), which show the field behavior within the bulk of the material. The curves have been calculated using Eq. (25) for $4c^{2}\kappa_{1}/\Omega^{2}=1000$ for $\kappa_{0}=2|\kappa_{1}|$ (period-doubling state) and $\kappa_{0}=2.5|\kappa_{1}|$ (non period-doubling state). Different colors refer to different times in units of the half period $T_{0}/2=\pi/\Omega$, as indicated in the colorbars. The black curves in both panels indicate the period-averaged magnetic field.
@@
$\frac{c^{2}}{4\pi\lambda_{L}^{2}(t)}=e_{\star}^{2}\langle\bar{b}^{2}(\mathbf{x},t)\rangle_{\mathrm{space}}\,$, (23)
@@
$\left\langle\frac{c^{2}}{4\pi\lambda_{L}^{2}(t)}\right\rangle_{\!\!\!\mathrm{period}}\!\!\!\!=e_{\star}^{2}b_{0}^{2}\langle\cos^{2}(\Omega t/2)\rangle_{\mathrm{period}}=\frac{e_{\star}^{2}b_{0}^{2}}{2},$ (24)
showing that the Meissner effect is expected to occur also in the period-doubling phases.
@@
with $\mathrm{ce}_{n}(y,\tau)$ time-periodic cosine-Mathieu functions of the first kind [MathieuNote], and the $\Phi_{n}(x)$ functions defined as
@@ VI Conclusion
For even values of $N$, we coupled the system to a background electromagnetic vector potential and demonstrated that all equilibrium and non-equilibrium phases with broken O($N$) symmetry exhibit a Meissner effect, in the sense that the electromagnetic field acquires a mass. Remarkably, when discrete time-translation symmetry is broken and an external magnetic field is applied, a portion of the magnetic flux can penetrate the system, forming a standing-wave pattern within the sample. We dubbed such pattern a Meissner polariton, a collective mode emerging from the interaction of order parameter oscillations with light. Furthermore, we found that increasing the drive strength can strongly enhance the imaginary part of the optical conductivity, even in the absence of O($N$) symmetry breaking or a Meissner effect.
Our results have broad implications. They establish that dynamical driving can induce emergent ordered phases, offering a unified theoretical framework for light-induced symmetry breaking phenomena observed not only in superconductors, but also in charge-density-wave, antiferromagnetic, and ferroelectric systems. Moreover, we showed that drive-induced symmetry-broken phases can display a time-dependent Meissner effect, providing a potential explanation for the recently reported magnetic field expulsion in optically driven YBCO [Cavalleri24]. We also demonstrated that near a non-equilibrium superconducting phase, the imaginary part of the optical conductivity can exhibit a pronounced $1/\omega$ scaling over several frequency decades, mimicking superconducting behavior even without a genuine Meissner response.
Promising directions for future work include applying our framework to microscopic models with explicit electronic degrees of freedom and extending it to more complex field theories that feature the competition or coexistence of multiple order parameters [Diessel_inprep].
@@ Acknowledgments
We gratefully acknowledge discussions A. Cavalleri, A. Chiocchetta, E. Demler, J.S. Dodge, P.E. Dolgirev, F. Glerean, M. Mitrano, and P.A. Volkov. This research was supported by NSF Grant DMR-2245246 and by the Simons Collaboration on Ultra-Quantum Matter which is a grant from the Simons Foundation (651440, S. S.). The Flatiron Institute is a division of the Simons Foundation. O.K.D acknowledges support from the NSF through a grant for ITAMP at Harvard University. P.M.B. acknowledges support by the German National Academy of Sciences Leopoldina through Grant No. LPDS 2023-06 and the Gordon and Betty Moore Foundation’s EPiQS Initiative Grant GBMF8683.
@@ Appendix A Hubbard-Stratonovich transformation and saddle point
In this section we present a derivation of the large-$N$ equations used in the main text. We start by writing the Lagrangian Eq. (2) on the Keldysh double time contour and we add a dissipative term [kamenev2023Book]. We also define $b_{\mathrm{cl}}$ and $b_{\mathrm{q}}$ as the classical and quantum components of $b$, respectively. We then perform a Hubbard-Stratonovich transformation to decouple the interaction term to get the Lagrangian density
@@ Appendix B Expressions for the coefficients $C_{n}(q)$ in the steady state
$\begin{split}\Big\{\partial_{t}^{3}+3\gamma\partial_{t}^{2}+\left[4q^{2}+4\lambda(t)+2\gamma^{2}\right]\partial_{t}\\ +\left[4\gamma q^{2}+4\gamma\lambda_{\pm}(t)+2\dot{\lambda}(t)\right]\Big\}C(q,t)=4\gamma T\,.\end{split}$ (40)
@@ Appendix C Iterative solution of Floquet equations
    Update $\lambda_{0}$ by imposing $\mathrm{det}\mathcal{K}(\omega)=0$ at $\omega=0$ or $\Omega/2$,
@@
    Calculate $\bar{b}_{n}$ up to a constant $\mathcal{B}$ from the kernel of $\mathcal{K}(\omega)$ ($\bar{b}_{n}=\mathcal{B}\,\hat{b}_{n}$ with $\sum_{n}|\hat{b}_{n}|^{2}=1$),
@@ Appendix D Determination of finite-$q$ instabilities
with $\tilde{\epsilon}_{\mathbf{q}}=\sqrt{q^{2}+\lambda_{0}-\gamma^{2}/4}$, and $\lambda_{\ell}$ are calculated numerically by solving the equations in real time and Fourier transforming $\lambda(t)$ in the steady state. The Floquet quasibands $E_{n}(\mathbf{q})$ are defined as the poles of $G^{R}_{nm}(\mathbf{q},\omega)$ for complex $\omega$. We can numerically calculate $E_{n}(\mathbf{q})$ by calculating the eigenvalues $\tilde{E}_{n}(\mathbf{q})$ of the matrix
@@
While each one of the blocks is an infinite matrix, we truncate them to $N\times N$ matrices to make numerical calculations amenable. In the phase diagram in Fig. 1 we used $N=11$. The Floquet quasibands are given by $E_{n}(\mathbf{q})=\tilde{E}_{n}(\mathbf{q})-i\gamma/2$. Note that $E_{n}(\mathbf{q})$ obeys $E_{n+m}(\mathbf{q})=E_{n}(\mathbf{q})+m\Omega$. To detect an instability, we focus on the imaginary part of $E_{n}(\mathbf{q})$. Due to causality, the poles of the retarded Green’s function should be located only in the lower complex half plane. Whenever the imaginary part of $E_{n}(\mathbf{q})$ turns zero for some $\mathbf{q}=\mathbf{Q}$ or becomes even positive, we say that the system has an instability. If this occurs at a nonzero value of $\mathbf{Q}$, then we have a finite momentum instability. To distinguish between period-preserving and period-doubling instabilities, we look at $\mathrm{Re}\,E_{n}(\mathbf{Q})]\,\mathrm{mod}\,\Omega$. If this quantity equals 0 or $\Omega$, then the instability is period-preserving, if instead it equals $\Omega/2$, then it is period-doubling. Note that these are the only possibilities. Typical behaviors of the imaginary parts of $E_{n}(\mathbf{q})$ are shown in the right column of Fig. 4, where a $\times2$ truncation was performed. Despite the small size of the truncated Green’s function, the qualitative behavior of $\mathrm{Im}\,E_{n}(\mathbf{q})$ is captured correctly.
@@ E.2 Meissner effect and optical conductivity
where $\eta=0$ for period-preserving solutions and $\eta=1$ for period doubling ones. Integrating the last term by parts, we can see that the period-averaged Meissner effect is given by the simple formula $\mathcal{K}_{n}^{\alpha\beta}(\omega\to 0)=2\sum_{m=-\infty}^{+\infty}\bar{b}_{\mathbf{Q},m}\bar{b}_{-\mathbf{Q},n-m+\eta}\,\delta_{\alpha\beta}$. Fourier transforming to real time and real space, we get a space-time-dependent London penetration depth given by

### v2

@@ III.1.4 Effect of finite temperature: Fluctuations shift the critical point
In the large frequency limit, $\Omega\gg\sqrt{|r_{1}|}$, $\gamma$, we can also approximate $C_{n}(q)\approx\frac{T}{q^{2}+\lambda_{0}}\delta_{n,0}$ and then calculate the critical value of $r_{0}$ for which the superconducting phase emerges. To this end, we use the definition of $\lambda_{0}$ as in Eq. (9) at $n=0$ and setting $b_{n}=0$. Then, we make use of the stability condition Eq. (11) (and replace $r_{0}$ with $\lambda_{0}$). The new critical value for $r_{0,c}$ is then given by:
@@ IV Full numerical solution
Finite-momentum instabilities— We also performed a stability analysis of each phase obtained by numerically solving Eqs. (4). We found that especially for larger values of $r_{1}/\Omega^{2}$ the normal phase is unstable to a finite-momentum symmetry broken state, where the order parameter displays oscillations as a function of time with a characteristic frequency of $\Omega$ (period-preserving state) of $\Omega/2$ (period-doubling state). For details of this analysis see App. D. As shown in App. E, such phases also exhibit a Meissner effect, which is why we dubbed them as pair density waves (PDW) in Fig. 1 (marked by red or blue hatches). Similarly to the analysis in Sec. III.1.3, the magnitude $Q=|\mathbf{Q}|$ of the momentum at which the instability occurs is approximately set by the condition $\tilde{\epsilon}_{\mathbf{Q}}=\sqrt{Q^{2}+\lambda_{0}-\gamma^{2}/4}=\Omega/2$. For a finite-momentum instability, the order parameter can take two different forms:
@@
with $f_{\nu}(t)$ a time-periodic function with period $(2\pi)/\nu$. In the first case the O($N$) symmetry is broken down to O(N1N-1) and both the translation and the spatial rotation symmetries are broken. In the second case, a subgroup of the combination of translations and O($N$) rotations remains preserved, while spatial rotations are broken and the global O($N$) symmetry is broken down to O(N2N-2). More complex patterns are in principle possible as one is allowed to construct any linear combination of wave vectors $\mathbf{q}$ satisfying $|\mathbf{q}|=|\mathbf{Q}|$. The determination of which finite-$q$ configuration is realized is beyond our scope and it would require solving a generalization of Eqs. (4) where $\bar{b}$ becomes spatially nonuniform. We also remark that our analysis (detailed in App. D) relies on the assumption that the finite-$q$ order parameter develops continuously upon crossing the transition line. If the transition happens to be first-order, the boundaries of the finite-momentum phases in Fig. 1 might change. We also found a regime (yellow region in Fig. 1), where $\lambda(t)$ is not a periodic function. Here, the correlation function $C_{\mathbf{q}}(t)$ has a large peak at finite $\mathbf{q}$, which is why we dubbed this phase a quasi-periodic PDW. Finite momentum instabilities in parametrically driven bosonic systems were discussed also in Refs. [11; 32; 33; 31].
@@ V Electromagnetic response
Expressions for the electromagnetic kernel in the steady state can be found in App. E.
@@ V.1 Meissner effect
Figure 5: Behavior of the magnetic field inside the material for a superconducting state with an order parameter that is either uniform or carries a finite momentum in the form of Eq. (17b). The $x$-axis coordinate indicates the distance from an interface to vacuum. In panels (a) and (b) we show a state with period doubling, whereas in panels (c) and (d) we have a superconducting state without period doubling, with an order parameter that oscillates around a finite offset. Panels (a) and (c) display the magnetic field profiles near the vacuum interface, in contrast to panels (b) and (d), which show the field behavior within the bulk of the material. The insets in panels (a) and (c) show the time series of the magnetic field at $x\Omega/(2c)=10$ for $\kappa_{0}=2|\kappa_{1}|$ and at $x\Omega/(2c)=0.05$ for $\kappa_{0}=2.5|\kappa_{1}|$, respectively. The curves have been calculated using Eq. (25) for $4c^{2}\kappa_{1}/\Omega^{2}=1000$ for $\kappa_{0}=2|\kappa_{1}|$ (period-doubling state) and $\kappa_{0}=2.5|\kappa_{1}|$ (non period-doubling state). Different colors refer to different times in units of the half period $T_{0}/2=\pi/\Omega$, as indicated in the colorbars. The black curves in both panels indicate the period-averaged magnetic field.
@@
$\frac{c^{2}}{4\pi\lambda_{L}^{2}(t)}=e_{\star}^{2}\langle|\bar{\phi}(\mathbf{x},t)|^{2}\rangle_{\mathrm{space}}\,$, (23)
@@
$\left\langle\frac{c^{2}}{4\pi\lambda_{L}^{2}(t)}\right\rangle_{\!\!\!\mathrm{period}}\!\!\!\!=e_{\star}^{2}|\phi_{0}|^{2}\langle\cos^{2}(\Omega t/2)\rangle_{\mathrm{period}}=\frac{e_{\star}^{2}|\phi_{0}|^{2}}{2},$ (24)
showing that the Meissner effect is expected to occur also in the period-doubling phases. Here $|\phi_{0}|^{2}=\langle|\bar{\phi}(\mathbf{x},t)|^{2}\rangle_{\text{space,time}}$.
@@
with $\mathrm{ce}_{n}(y,\tau)$ time-periodic cosine-Mathieu functions of the first kind [1], and the $\Phi_{n}(x)$ functions defined as
@@
Finally, we note that in both the same-period and period-doubled phases (Fig. 5, insets of panels (a) and (c)), the magnetic field $\bm{B}(\bm{x},t)$ oscillates with period $T_{0}$. This follows from the fact that the inverse squared penetration depth is proportional to $|\phi(t)|^{2}$, see Eq. (23). While $\phi(t)$ can be $T_{0}$- or $2T_{0}$-periodic depending on whether $\nu=0$ or $\nu=1/2$ in Eq. (6), its square $|\phi(t)|^{2}$ is always $T_{0}$-periodic. In fact, if it is $T_{0}$ periodic, it will always contain a time-independent component (see Eq. (6)), whereas this vanishes in period-doubling phases. Consequently, $\bm{B}(\bm{x},t)$ inherits the fundamental period $T_{0}$ in both phases. We also note that, in the theory in Eq. (18), all gauge invariant quantities must contain $\phi(t)$ as $|\phi(t)|^{2}$, making it impossible to observe $2T_{0}$ oscillations. Conversely, if the theory in Eq. (2) is not coupled to a gauge field and if the order parameter $b(t)$ itself is an observable, $2T_{0}$ oscillations are in principle possible to detect.
@@ VI Conclusion
[…]
Our results have broad implications. They establish that dynamical driving can induce emergent ordered phases, offering a unified theoretical framework for light-induced symmetry breaking phenomena observed not only in superconductors, but also in charge-density-wave, antiferromagnetic, and ferroelectric systems. Moreover, we showed that drive-induced symmetry-broken phases can display a time-dependent Meissner effect, providing a potential explanation for the recently reported magnetic field expulsion in optically driven YBCO [19]. We also demonstrated that near a non-equilibrium superconducting phase, the imaginary part of the optical conductivity can exhibit a pronounced $1/\omega$ scaling over several frequency decades, mimicking superconducting behavior even without a genuine Meissner response.
Promising directions for future work include applying our framework to microscopic models with explicit electronic degrees of freedom, and extending it to more complex field theories that feature the competition or coexistence of multiple order parameters [15].
@@ Acknowledgments
We gratefully acknowledge discussions A. Cavalleri, A. Chiocchetta, E. Demler, J.S. Dodge, P.E. Dolgirev, F. Glerean, M. Mitrano, P.A. Volkov, and C.P. Zelle. This research was supported by NSF Grant DMR-2245246 and by the Simons Collaboration on Ultra-Quantum Matter which is a grant from the Simons Foundation (651440, S. S.). The Flatiron Institute is a division of the Simons Foundation. O.K.D acknowledges support from the NSF through a grant for ITAMP at Harvard University. P.M.B. acknowledges support by the German National Academy of Sciences Leopoldina through Grant No. LPDS 2023-06 and the Gordon and Betty Moore Foundation’s EPiQS Initiative Grant GBMF8683.
@@ Appendix A Hubbard-Stratonovich transformation and saddle point
In this section we present a derivation of the large-$N$ equations used in the main text. We start by writing the Lagrangian Eq. (2) on the Keldysh double time contour and we add a dissipative term [29; 57; 58]. We also define $b_{\mathrm{cl}}$ and $b_{\mathrm{q}}$ as the classical and quantum components of $b$, respectively. We then perform a Hubbard-Stratonovich transformation to decouple the interaction term to get the Lagrangian density
@@ Appendix B Expressions for the coefficients $C_{n}(q)$ in the steady state
$\begin{split}\Big\{&\partial_{t}^{3}+3\gamma\partial_{t}^{2}+\left[4q^{2}+4\lambda(t)+2\gamma^{2}\right]\partial_{t}\\ &+\left[4\gamma q^{2}+4\gamma\lambda_{\pm}(t)+2\dot{\lambda}(t)\right]\Big\}C(q,t)=4\gamma T\,.\end{split}$ (40)
@@ Appendix C Iterative solution of Floquet equations
    Update $\lambda_{0}$ by imposing $\mathrm{det}\mathcal{K}(\nu)=0$ at $\omega=0$ or $\Omega/2$,
@@
    Calculate $\bar{b}_{n}$ up to a constant $\mathcal{B}$ from the kernel of $\mathcal{K}(\nu)$ ($\bar{b}_{n}=\mathcal{B}\,\hat{b}_{n}$ with $\sum_{n}|\hat{b}_{n}|^{2}=1$),
@@ Appendix D Determination of finite-$q$ instabilities
with $\tilde{\epsilon}_{\mathbf{q}}=\sqrt{q^{2}+\lambda_{0}-\gamma^{2}/4}$, and $\lambda_{\ell}$ are calculated numerically by solving the equations in real time and Fourier transforming $\lambda(t)$ in the steady state. The Floquet quasi-bands $E_{n}(\mathbf{q})$ are defined as the poles of $G^{R}_{nm}(\mathbf{q},\omega)$ for complex $\omega$. We can numerically calculate $E_{n}(\mathbf{q})$ by calculating the eigenvalues $\tilde{E}_{n}(\mathbf{q})$ of the matrix
@@
While each one of the blocks is an infinite matrix, we truncate them to $N\times N$ matrices to make numerical calculations amenable. In the phase diagram in Fig. 1 we used $N=11$. The Floquet quasi-bands are given by $E_{n}(\mathbf{q})=\tilde{E}_{n}(\mathbf{q})-i\gamma/2$. Note that $E_{n}(\mathbf{q})$ obeys $E_{n+m}(\mathbf{q})=E_{n}(\mathbf{q})+m\Omega$. To detect an instability, we focus on the imaginary part of $E_{n}(\mathbf{q})$. Due to causality, the poles of the retarded Green’s function should be located only in the lower complex half plane. Whenever the imaginary part of $E_{n}(\mathbf{q})$ turns zero for some $\mathbf{q}=\mathbf{Q}$ or becomes even positive, we say that the system has an instability. If this occurs at a nonzero value of $\mathbf{Q}$, then we have a finite momentum instability. To distinguish between period-preserving and period-doubling instabilities, we look at $\mathrm{Re}\,E_{n}(\mathbf{Q})]\,\mathrm{mod}\,\Omega$. If this quantity equals 0 or $\Omega$, then the instability is period-preserving, if instead it equals $\Omega/2$, then it is period-doubling. Note that these are the only possibilities. Typical behaviors of the imaginary parts of $E_{n}(\mathbf{q})$ are shown in the right column of Fig. 4, where a $\times2$ truncation was performed. Despite the small size of the truncated Green’s function, the qualitative behavior of $\mathrm{Im}\,E_{n}(\mathbf{q})$ is captured correctly.
@@ E.2 Meissner effect and optical conductivity
where $\eta=0$ for period-preserving solutions and $\eta=1$ for period doubling ones. Integrating the last term by parts, we get
$\int_{\mathbf{q}}\int_{-\infty}^{+\infty}\!\frac{d\nu}{2\pi}\,\bm{G}^{K}_{0,n}(\nu)\,\delta_{\alpha\beta}=-\int_{\mathbf{q}}\int_{-\infty}^{+\infty}\!\frac{d\nu}{2\pi}\,\partial_{q_{\alpha}}\bm{G}^{K}_{0,n}(\nu)\,q_{\beta}\,.$ (53)
Using Eqs. (50a) and (50b), one obtains
$\partial_{q_{\alpha}}\bm{G}^{K}(\nu)=(2q_{\alpha})\sum_{m=-\infty}^{+\infty}\left[\bm{G}^{R}_{0,m}(\nu)\bm{G}^{K}_{m,n}(\nu)+\bm{G}^{K}_{0,m}(\nu)\bm{G}^{A}_{m,n}\right]\,.$ (54)
Plugging (53) and (54) into Eq. (52), we obtain the simple formula $\mathcal{K}_{n}^{\alpha\beta}(\omega\to 0)=2\sum_{m=-\infty}^{+\infty}\bar{b}_{\mathbf{Q},m}\bar{b}_{-\mathbf{Q},n-m+\eta}\,\delta_{\alpha\beta}$. Fourier transforming to real time and real space, we get a space-time-dependent London penetration depth given by